# Delving into Temperature Scaling for Adaptive Conformal Prediction

### v2

@@ 7 Conclusion
In this paper, we introduce Conformal Temperature Scaling (ConfTS), a modification to Temperature Scaling that enhances the efficiency of adaptive conformal prediction. ConfTS optimizes the temperature value by minimizing the efficiency gap on a held-out validation set. The obtained temperature would encourage the prediction sets to approximate optimal sets of high efficiency while maintaining the marginal coverage. Extensive experiments show that with ConfTS, the prediction set can be efficient and have better conditional coverage performance. By enhancing adaptive conformal prediction in a post-hoc manner, ConfTS can be easily implemented within any deep learning framework without sacrificing predictive performance. We hope that the insights from this study can serve as a guideline for researchers to effectively incorporate temperature scaling into conformal prediction.
Limitations.
Our proposed method is effective only under the assumption of data exchangeability. Future research could explore the development of post-hoc rescaling techniques to enhance conformal prediction performance under various distribution shifts, such as covariance shift and semantic shift.
Appendix A Conformal prediction methods and metrics
@@
where $\mathds{1}$ is the indicator function and $\mathcal{D}_{test}$ denotes the test dataset. The coverage rate measures the percentage of samples whose prediction set contains the true label, i.e., an empirical estimation for $\mathbb{P}(Y\in\mathcal{C}(X))$. Average size measures the efficiency of prediction sets. The prediction sets should both provide valid coverage (defined in Eq. (2)) and efficiency (i.e., small prediction sets). Smaller prediction sets are often preferred since they are more informative in practice (Vovk, 2012; Angelopoulos et al., 2021b).
Moreover, we use size-stratified coverage violation (SSCV) (Angelopoulos et al., 2021b) to measure the performance of prediction sets on conditional coverage. Specifically, considering a disjoint set-size strata $\{S_{i}\}_{i=1}^{N_{s}}$, where $\bigcup_{i=1}^{N_{s}}S_{i}=\{1,2,\cdots,|\mathcal{Y}|\}$. Then, we define the indexes of examples stratified by the prediction set size by $\mathcal{J}_{j}=\{i:|\mathcal{C}(\bm{x}_{i})|\in S_{j}\}$. Formally, we can define the SSCV as:
$\text{SSCV}=\sup_{j}\left|\frac{\left|\{i\in\mathcal{J}_{j}:y_{i}\in\mathcal{C}\left(\bm{x}_{i}\right)\}\right|}{\left|\mathcal{J}_{j}\right|}-(1-\alpha)\right|.$ (12)
This metric measures the maximum deviation from the target coverage rate $1-\alpha$ across all strata. A lower SSCV value generally indicates better conditional coverage performance of the prediction sets.
Appendix B Confidence calibration methods
@@
Appendix C Experiment Setup of Section 3
Datasets. We use two datasets in our study: ImageNet (Deng et al., 2009) and CIFAR-100 (Krizhevsky et al., 2009). On ImageNet, we split the test dataset including 50,000 images into 10,000 images for the calibration set and 40,000 images for the test set; on CIFAR-100, we split the test dataset including 10,000 images into 4,000 images for the calibration set and 6,000 for the test set. Then, we split the calibration set into two subsets of equal size: one is the validation set used for confidence calibration, while the other half is the conformal set used for conformal calibration.
Models. We employ three pre-trained classifiers: ResNet18, ResNet50, ResNet101 (He et al., 2016) from TorchVision (Paszke et al., 2019); on CIFAR-100, we train models from scratch. The models are trained for 100 epochs using SGD with a momentum of 0.9, a weight decay of 0.0005, and a batch size of 258. We set the initial learning rate as 0.1 and reduce it by a factor of 5 at 60 epochs. For ConfTr, we set the smoothing parameter $T=0.1$, penalty term $\kappa=1$, and hyperparameter $\lambda=1$, using the same setups. We conduct all the experiments on NVIDIA GeForce RTX 4090.
Conformal prediction algorithms. We leverage APS and RAPS to generate prediction sets at an error rate $\alpha=0.1$, and the hyperparameters are set to be $k_{reg}=1$ and $\lambda=0.001$.
Evaluation. For the evaluation metrics, we employ coverage and average size (see Appendix A) to evaluate the performance of prediction sets and utilize ECE to measure the miscalibration (see Section 3.1). All experiments are repeated 20 times with different seeds, and we report the average performance.
Appendix D Results of training-time calibration methods
In this section, we report the results of how training-time calibration methods and Bayesian deep learning with Laplace approximation affect the conformal prediction methods. Specifically, we employ label smoothing and mixup to train a ResNet50 model on the CIFAR100 dataset from scratch and utilize Laplace approximation in a post-hoc manner. For label smoothing, we set the hyperparameter $\alpha=0.05$, and for mixup, the hyperparameter is set to be $\alpha=0.1$. The training details are available in Appendix C. In Table 5, we show that employing these calibration methods enlarges the prediction sets of APS and RAPS, which is consistent with our results in the main paragraph. For example, with label smoothing, the average size of APS increases from $4.91$ to $11.9$.
Table 5: Results of different calibration methods using ResNet50 on CIFAR-100. $\downarrow$” indicates smaller values are better. “” and “” indicate whether the performance is superior/inferior to the baseline. Bold numbers are superior results. Results show that existing training-time calibration methods and Bayesian deep learning often hurt the efficiency of adaptive conformal prediction.
$\alpha=0.1$ $\alpha=0.05$
Method Baseline LabelSmoothing Mixup Bayesian Baseline LabelSmoothing Mixup Bayesian
Accuracy 0.77 0.78 0.78 0.77 0.77 0.78 0.78 0.77
ECE(%) $\downarrow$ 8.79 4.39 2.96 4.3 8.79 4.39 2.96 4.3
Avg.size(APS) $\downarrow$ 4.91 11.9 12.5 7.55 11.1 19.8 20.1 15.6
Coverage(APS) 0.90 0.90 0.90 0.90 0.95 0.95 0.95 0.95
Avg.size(RAPS) $\downarrow$ 2.56 9.5 10.2 6.46 6.95 14.5 15.5 9.34
Coverage(RAPS) 0.90 0.90 0.90 0.90 0.95 0.95 0.95 0.95
@@ Appendix D Results of training-time calibration methods
Appendix E Why numerical error occurs under an exceedingly small temperature?
@@
Appendix F Proofs
F.1 Proof for Proposition 3.1
We start by showing several lemmas: the Lemma F.1, Lemma F.2 and Lemma F.3.
Lemma F.1.
@@ Appendix F Proofs
We continue by showing the Eq. (19) at $k=K-1$. The Lemma F.1 gives us that
@@
Considering the Lemma F.2, the Eq. (24) implies that
@@
Proposition F.4 (Restatement of Proposition 3.1).
@@
Since the scaling parameter $t_{0}$ does not change the order of $(\tilde{f}_{1},\tilde{f}_{2},\cdots,\tilde{f}_{K})$, i.e. $\tilde{f}_{1}>\tilde{f}_{2}>\cdots>\tilde{f}_{K}$ and $\alpha\in(0,1)$, then by the Lemma F.3, we have $\mathcal{S}(\bm{x},y,t_{0})<\mathcal{S}(\bm{x},y,t)$. ∎
F.2 Proof for Corollary 3.2
Corollary F.5 (Restatement of Corollary 3.2).
@@
Theorem F.7.
Under the assumption (26), there exists constants $c_{1},\gamma\in(0,1]$ such that
@@
Continuing from Lemma F.6, we can get
@@
Appendix G Results of ConfTS on CIFAR-100
@@
Table 6: Performance comparison of the baseline and ConfTS on CIFAR-100 dataset. We employ five models trained on CIFAR-100. “*” denotes significant improvement (two-sample t-test at a 0.1 confidence level). $\downarrow$” indicates smaller values are better. Bold numbers are superior results. Results show that our ConfTS can improve the performance of APS and RAPS, maintaining the desired coverage rate.
@@
Appendix H Results of ConfTS on ImageNet-V2
Table 7: Performance comparison of conformal prediction with baseline and ConfTS under distribution shifts. “*” denotes significant improvement (two-sample t-test at a 0.1 confidence level). $\downarrow$” indicates smaller values are better. Bold numbers are superior results. Results show that ConfTS can improve the efficiency of APS and RAPS on a new distribution.
Metrics ResNet50 DenseNet121 VGG16 ViT
Baseline ConfTS Baseline ConfTS Baseline ConfTS Baseline ConfTS
Avg.size(APS) $\downarrow$ 24.6 11.9* 50.3 13.3* 27.2 17.9* 34.2 10.1*
Coverage(APS) 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
Coverage(RAPS) 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
Appendix I Results of ConfTS on RAPS with various penalty terms
Recall that the RAPS method modifies APS by including a penalty term $\lambda$ (see Eq. (6)). In this section, we investigate the performance of ConfTS on RAPS with various penalty terms. In particular, we employ the same model architectures with the main experiment on ImageNet (see Section 5.1) and generate prediction sets with RAPS ($k_{reg}=1$) at an error rate $\alpha=0.1$, varying the penalty $\lambda\in\{0.002,0.004,0.006,0.01,0.015,0.02\}$ and setting $k_{reg}$ to 1. Table 8 and 9 show that our ConfTS can enhance the efficiency of RAPS across various penalty values.
Table 8: Performance of ConfTS on RAPS with various penalty terms $\lambda\in\{0.002,0.004,0.006\}$ at ImageNet. “*” denotes significant improvement (two-sample t-test at a 0.1 confidence level). $\downarrow$” indicates smaller values are better. Bold numbers are superior results. Results show that our ConfTS can enhance the efficiency of RAPS across various penalty values.
@@ Appendix I Results of ConfTS on RAPS with various penalty terms
Table 9: Performance of ConfTS on RAPS with various penalty terms $\lambda\in\{0.01,0.015,0.02\}$ at ImageNet. “*” denotes significant improvement (two-sample t-test at a 0.1 confidence level). $\downarrow$” indicates smaller values are better. Bold numbers are superior results. Results show that our ConfTS can enhance the efficiency of RAPS across various penalty values.
@@
Appendix J Results of ConfTS on SAPS
@@
Table 10: Performance of ConfTS on SAPS with various penalty terms $\lambda\in[0.005,0.01,0.02]$. “*” denotes significant improvement (two-sample t-test at a 0.1 confidence level). $\downarrow$” indicates smaller values are better. Bold numbers are superior results. Results show that our ConfTS can enhance the efficiency of SAPS across various penalty values.
@@
Table 11: Performance of ConfTS on SAPS with various penalty terms $\lambda\in\{0.03,0.05,0.1\}$. “*” denotes significant improvement (two-sample t-test at a 0.1 confidence level). $\downarrow$” indicates smaller values are better. Bold numbers are superior results. Results show that our ConfTS can enhance the efficiency of SAPS across various penalty values.

### v3

@@ 5 Conclusion
In this paper, we investigate the relationship between two uncertainty estimation frameworks: confidence calibration and conformal prediction. We make two discoveries about this relationship: firstly, existing confidence calibration methods would lead to larger prediction sets for adaptive conformal prediction; secondly, high-confidence prediction could enhance the efficiency of adaptive conformal prediction. We prove that applying a smaller temperature to a prediction could lead to more efficient prediction sets on expectation. Inspired by this, we propose a variant of temperature scaling, Conformal Temperature Scaling (ConfTS), which rectifies the optimization objective toward generating efficient prediction sets. Our method can be extended to other post-hoc calibrators for improving conformal predictors. Extensive experiments demonstrate that our method can enhance existing adaptive conformal prediction methods, in both image and text classification tasks. Our work challenges the conventional wisdom of utilizing confidence calibration for conformal prediction, and we hope it can inspire specially-designed methods to improve the two frameworks of uncertainty estimation.
Limitations. In this work, the conclusions of our analysis are mostly for adaptive conformal prediction methods, without generalizing to the LAC score. In addition, the proposed method only focuses on enhancing the efficiency of prediction sets and may not help in conditional coverage, similar to current training methods for conformal prediction. We believe it can be interesting to design loss functions specifically tailored for improving conditional coverage, in future works.
Appendix A Related Work
Conformal prediction. Conformal prediction (Papadopoulos et al., 2002; Vovk et al., 2005) is a statistical framework for uncertainty qualification. Previous works have applied conformal prediction across various domains, including regression (Lei & Wasserman, 2014; Romano et al., 2019), image classification (Angelopoulos et al., 2021b; Huang et al., 2024), hyperspectral image classification (Liu et al., 2024), object detection (Angelopoulos et al., 2021a; Teng et al., ), outlier detection (Bates et al., 2023; Guan & Tibshirani, 2022; Liang et al., 2022), and large language models (Kumar et al., 2023). Conformal prediction is also deployed in other real-world applications, such as human-in-the-loop decision-making (Straitouri et al., 2023; Cresswell et al., 2024), automated vehicles (Bang et al., 2024), and scientific computation (Moya et al., 2024).
Some methods leverage post-hoc techniques to enhance prediction sets (Romano et al., 2020; Angelopoulos et al., 2021b; Ghosh et al., 2023; Huang et al., 2024). For example, Adaptive Prediction Sets (APS) (Romano et al., 2020) calculates the score by accumulating the sorted softmax values in descending order. However, the softmax probabilities typically exhibit a long-tailed distribution, and thus those tail classes are often included in the prediction sets. To alleviate this issue, Regularized Adaptive Prediction Sets (RAPS) (Angelopoulos et al., 2021b) exclude tail classes by appending a penalty to these classes, resulting in efficient prediction sets. These post-hoc methods often employ temperature scaling for better calibration performance (Angelopoulos et al., 2021b; Lu et al., 2022; Gibbs et al., 2023; Lu et al., 2023). In our work, we show that existing confidence calibration methods harm the efficiency of adaptive conformal prediction.
Some works propose training time regularizations to improve the efficiency of conformal prediction (Colombo & Vovk, 2020; Stutz et al., 2022; Einbinder et al., 2022; Bai et al., ; Correia et al., 2024). For example, uncertainty-aware conformal loss function (Einbinder et al., 2022) optimizes the efficiency of prediction sets by encouraging the non-conformity scores to follow a uniform distribution. Moreover, conformal training (Stutz et al., 2022) constructs efficient prediction sets by prompting the threshold to be less than the non-conformity scores. In addition, information-based conformal training (Correia et al., 2024) incorporates side information into the construction of prediction sets. In this work, we focus on post-hoc training methods, which only require the pre-trained models for conformal prediction. ConfTS is easy to implement and requires low computation resources.
Confidence calibration.
Confidence calibration has been studied in various contexts in recent years. Some works address the miscalibration problem by post-hoc methods, including histogram binning (Zadrozny & Elkan, 2001) and Platt scaling (Platt et al., 1999). Besides, regularization methods like entropy regularization (Pereyra et al., 2017) and focal loss (Mukhoti et al., 2020) are also proposed to improve the calibration performance of deep neural networks. In addition, data augmentation techniques also play a role in confidence calibration, including label smoothing (Müller et al., 2019) and mixup (Thulasidasan et al., 2019). A concurrent work (Dabah & Tirer, 2024) also investigates the effects of temperature scaling on conformal prediction. However, they only focus on the temperature scaling and do not extend the conclusion to other post-hoc and training methods of confidence calibration. In this work, we provide a more comprehensive analysis with both post-hoc and training methods of confidence calibration. In addition to the analysis, we also provide a theoretical explanation and introduce a novel method to automatically optimize the parameters of post-hoc calibrators.
Appendix B Conformal prediction metrics
@@
where $\mathds{1}(\cdot)$ is the indicator function and $\mathcal{D}_{test}$ denotes the test dataset. The coverage rate measures the percentage of samples whose prediction set contains the true label, i.e., an empirical estimation for $\mathbb{P}\{Y\in\mathcal{C}(X)\}$. The average size measures the efficiency of prediction sets. For informative predictions (Vovk, 2012; Angelopoulos et al., 2021b), the prediction sets are preferred to be efficient (i.e., small prediction sets) while satisfying the valid coverage (defined in Eq. (2)).
Appendix C Confidence calibration methods
@@
Appendix D Experimental setups for motivation experiments
We conduct the experiments on CIFAR-100 (Krizhevsky et al., 2009). We split the test dataset including 10,000 images into 4,000 images for the calibration set and 6,000 for the test set. Then, we split the calibration set into two subsets of equal size: one is the validation set used for confidence calibration, while the other half is the conformal set used for conformal calibration. We train a ResNet50 model from scratch. For post-hoc methods, we train the model using standard cross-entropy loss, while for training methods, we use their corresponding specific loss functions. The training detail is presented in Section 4.1. We leverage APS and RAPS to generate prediction sets at an error rate $\alpha=0.1$, and the hyperparameters are set to be $k_{reg}=1$ and $\lambda=0.001$.
Appendix E Experiment results of LAC
E.1 How does confidence calibration affects LAC?
In this part, we investigate the connection between and confidence calibration methods. We employ three pre-trained classifiers: ResNet18, ResNet101 (He et al., 2016), DenseNet121 (Huang et al., 2017) on CIFAR-100, generating LAC prediction sets with $\alpha=0.1$. In Table 6, the results show that different post-hoc methods have varying impacts on LAC prediction sets, while all of them can maintain the desired coverage rate. For example, the original average size of ResNet18 with respect to THR is 2.23, increases to 2.40 with vector scaling, 2.34 with temperature scaling, and decreases to 2.20 with Platt scaling.
Table 6: The performance of LAC prediction sets employed with different calibration methods: baseline (BS), vector scaling (VS), Platt scaling (PS), and temperature scaling (TS). $\downarrow$ indicates smaller values are better.
Model Metrics BS VS PS TS
ResNet18 Avg.size $\downarrow$ 2.23 2.42 2.26 2.34
Coverage 0.90 0.90 0.90 0.90
ResNet101 Avg.size $\downarrow$ 1.88 1.98 1.83 1.83
Coverage 0.90 0.90 0.90 0.90
DenseNet121 Avg.size $\downarrow$ 1.68 1.68 1.69 1.65
Coverage 0.90 0.90 0.90 0.90
E.2 LAC with high-confidence prediction
In Figure 2, we compare the performance of LAC prediction sets deployed with different temperatures. We observe that when used with a small temperature, models tend to generate large prediction sets, while the coverage rate stabilizes at about 0.9, maintaining the marginal coverage. Moreover, we observe that models typically construct the smallest prediction set when the temperature approximates 1. Therefore, we cannot search for an appropriate temperature that benefits LAC.
(a) Coverage
(b) LAC
Figure 2: The performance comparison of prediction sets with different temperatures.
@@ E.2 LAC with high-confidence prediction
Appendix F Why numerical error occurs under an exceedingly small temperature?
@@
Appendix G Proofs
G.1 Proof for Proposition 3.1
We start by showing several lemmas: the Lemma G.1, Lemma G.2 and Lemma G.3.
Lemma G.1.
@@ Appendix G Proofs
We continue by showing the Eq. (18) at $k=K-1$. The Lemma G.1 gives us that
@@
Considering the Lemma G.2, the Eq. (23) implies that
@@
Proposition G.4 (Restatement of Proposition 3.1).
@@
Since the scaling parameter $t_{0}$ does not change the order of $(\tilde{f}_{1},\tilde{f}_{2},\cdots,\tilde{f}_{K})$, i.e. $\tilde{f}_{1}>\tilde{f}_{2}>\cdots>\tilde{f}_{K}$ and $\alpha\in(0,1)$, then by the Lemma G.3, we have $\mathcal{S}(\bm{x},y,t_{0})<\mathcal{S}(\bm{x},y,t)$. ∎
G.2 Proof for Corollary 3.2
Corollary G.5 (Restatement of Corollary 3.2).
@@
Theorem G.7.
Under the assumption equation 25, there exists constants $c_{1},\gamma\in(0,1]$ such that
@@
Continuing from Lemma G.6, we can get
@@
Appendix H Results of ConfTS on ImageNet-V2
Table 7: The performance comparison of conformal prediction with baseline and ConfTS under distribution shifts. “*” denotes significant improvement (two-sample t-test at a 0.1 confidence level). $\downarrow$” indicates smaller values are better. Bold numbers are superior results. Results show that ConfTS can improve the efficiency of APS and RAPS on a new distribution.
Metrics ResNet50 DenseNet121 VGG16 ViT
Baseline ConfTS Baseline ConfTS Baseline ConfTS Baseline ConfTS
Avg.size(APS) $\downarrow$ 24.6 11.9* 50.3 13.3* 27.2 17.9* 34.2 10.1*
Coverage(APS) 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
Coverage(RAPS) 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
Appendix I Results of ConfTS on CIFAR-100
@@
Table 8: The performance comparison of the baseline and ConfTS on CIFAR-100 dataset. We employ five models trained on CIFAR-100. “*” denotes significant improvement (two-sample t-test at a 0.1 confidence level). $\downarrow$” indicates smaller values are better. Bold numbers are superior results. Results show that our ConfTS can improve the performance of APS and RAPS, maintaining the desired coverage rate.
@@
Appendix J Results of ConfTS on RAPS with various penalty terms
Recall that the RAPS method modifies APS by including a penalty term $\lambda$ (see Eq. (6)). In this section, we investigate the performance of ConfTS on RAPS with various penalty terms. In particular, we employ the same model architectures with the main experiment on ImageNet (see Section 4.1) and generate prediction sets with RAPS ($k_{reg}=1$) at an error rate $\alpha=0.1$, varying the penalty $\lambda\in\{0.002,0.004,0.006,0.01,0.015,0.02\}$ and setting $k_{reg}$ to 1. Table 9 and 10 show that our ConfTS can enhance the efficiency of RAPS across various penalty values.
Table 9: The performance of ConfTS on RAPS with various penalty terms $\lambda\in\{0.002,0.004,0.006\}$ at ImageNet. “*” denotes significant improvement (two-sample t-test at a 0.1 confidence level). $\downarrow$” indicates smaller values are better. Bold numbers are superior results. Results show that our ConfTS can enhance the efficiency of RAPS across various penalty values.
@@
Table 10: The performance of ConfTS on RAPS with various penalty terms $\lambda\in\{0.01,0.015,0.02\}$ at ImageNet. “*” denotes significant improvement (two-sample t-test at a 0.1 confidence level). $\downarrow$” indicates smaller values are better. Bold numbers are superior results. Results show that our ConfTS can enhance the efficiency of RAPS across various penalty values.
@@
Appendix K Results of ConfTS on SAPS
@@
Table 11: The Performance of ConfTS on SAPS with various penalty terms $\lambda\in[0.005,0.01,0.02]$. “*” denotes significant improvement (two-sample t-test at a 0.1 confidence level). $\downarrow$” indicates smaller values are better. Bold numbers are superior results. Results show that our ConfTS can enhance the efficiency of SAPS across various penalty values.
@@
Table 12: The performance of ConfTS on SAPS with various penalty terms $\lambda\in\{0.03,0.05,0.1\}$. “*” denotes significant improvement (two-sample t-test at a 0.1 confidence level). $\downarrow$” indicates smaller values are better. Bold numbers are superior results. Results show that our ConfTS can enhance the efficiency of SAPS across various penalty values.
@@
(a) ResNet18
(b) ResNet50
Figure 4: (a)&(b): Average sizes of examples with different difficulties using APS on ResNet18 and ResNet50 respectively. Results show that ConfTS can maintain adaptiveness. (c)&(d) Average sizes of APS employed with ConfTS under various sizes of (c) conformal dataset (d) validation dataset. Results show that our ConfTS is robust to variations in the validation and conformal dataset size.
Appendix L ConfTS maintains the adaptiveness.
Adaptiveness (Romano et al., 2020; Angelopoulos et al., 2021b; Seedat et al., 2023) requires prediction sets to communicate instance-wise uncertainty: easy examples should obtain smaller sets than hard ones. In this part, we examine the impact of ConfTS on the adaptiveness of prediction sets and measure the instance difficulty by the order of the ground truth $o(y,\pi(\bm{x}))$. Specifically, we partition the sample by label order: 1, 2-3, 4-6, 7-10, 11-100, 101-1000, following (Angelopoulos et al., 2021b). Figure 4(a) and Figure 4(b) show that prediction sets, when applied with ConfTS, satisfy the adaptiveness property. Notably, employing ConfTS can promote smaller prediction sets for all examples ranging from easy to hard. Overall, the results demonstrate that APS with ConfTS succeeds in producing adaptive prediction sets: examples with lower difficulty obtain smaller prediction sets on average.
Appendix M Ablation study on the size of validation and calibration set.
In the experiment, ConfTS splits the calibration data into two subsets: validation set for tuning the temperature and conformal set for conformal calibration. In this part, we analyze the impact of this split on the performance of ConfTS by varying the validation and conformal dataset sizes from 3,000 to 8,000 samples while maintaining the other part at 5,000 samples. We use ResNet18 and ResNet50 on ImageNet, with APS at $\alpha=0.1$. Figure 4(c) and 4(d) show that the performance of ConfTS remains consistent across different conformal dataset sizes and validation dataset sizes. Based on these results, we choose a calibration set including 10000 samples and split it into two equal subsets for the validation and conformal set. In summary, the performance of ConfTS is robust to variations in the validation dataset and conformal dataset size.